# Exceptionally High Carrier Mobility in Hexagonal Diamond

### v1

@@ Abstract
Hexagonal diamond (h-diamond), or Lonsdaleite, has been reported to be a wide-bandgap semiconductor with high thermal conductivity and hardness. Our ab initio calculations show that h-diamond has exceptionally high carrier mobility. Along and directions, the hole mobilities at 300 K are 5631 and 5552 cm2V-1s-1, and the room-temperature electron mobilities are 11462 and 28464 cm2V-1s-1, respectively. These values are superior to the mobility of most known semiconductors including cubic diamond (c-diamond). The small effective masses in h-diamond, comparable to those in c-diamond, cannot explain the mobility difference between the two phases. For holes, scattering induced by transverse acoustic phonons is the predominant mechanism near room temperature in c-diamond, whereas considerably suppressed in d-diamond by selection rules. The high electron mobility in h-diamond can be attributed to the wavefunction at the conduction band minimum, which is extended and distributed primarily in the lattice interstitial, leading to weak coupling with scattering potentials. The temperature dependence of h-diamond is investigated as well, which deviates from the power-law relationship due to the significantly increased occupation of optical modes at elevated temperatures. Consequently, our findings reveal h-diamond as a promising high-mobility semiconductor, and elucidate the microscopic origin in terms of the carrier-phonon scattering mechanisms beyond conventional understandings based on simple parameters such as effective mass.
@@
Diamond, long prized as a gemstone, is also regarded as the “ultimate” wide-bandgap semiconductor material because of its excellent physical properties [21]. While most semiconductors may show good performance in certain properties, they often face inherent limitations in other critical characteristics. For example, GaAs has high electron mobility yet low hole mobility and thermal conductivity [2]; -Ga2O3 has an ultrawide bandgap and large breakdown field, whereas its carrier mobility and thermal conductivity are unsatisfactory [8]. In contrast, diamond offers a unique combination of a series of key properties, including its high carrier mobility, high thermal conductivity, ultra-wide bandgap, and high breakdown field [21]. These intrinsic characteristics position it at the frontier of future electronics operating under the most extreme conditions of high power, frequency, and temperature [18].
@@
In addition to the well-known cubic structure, diamond can also crystallize in the hexagonal phase. Hexagonal diamond (h-diamond), also referred to as Lonsdaleite [6], was both discovered in meteorites [9] and experimentally synthesized [3] in the 1960s. In contrast to cubic diamond (c-diamond) with a stacking sequence of ABCABC… along the crystallographic direction, it exhibits the ABAB… stacking. Although far less explored than the well-known cubic phase, existing studies have shown that the properties of h-diamond are comparable to, or even superior to, those of c-diamond. For example, Pan et al. [15] have revealed a 58% higher indentation strength of h-diamond than that of its cubic counterpart. Regarding the lattice thermal conductivity, studies on h-diamond report scattered values [25, 17, 4], which are approximately 20–30% lower than that of c-diamond, yet still considerably outperform most of the known materials.
@@
In this work, motivated by the recent success in the experimental synthesis of large-scale samples [22], we thoroughly investigate the carrier mobility of h-diamond, which is a key property that affects charge transport in semiconductors. Based on fully ab initio calculations, our results reveal that its electron mobility significantly exceeds that of most known semiconductor materials. We also shed light on the physical origin of its high carrier mobilities for both electrons and holes, which provides useful guidance for the discovery and design of novel high-mobility materials.
@@
The calculated carrier mobilities of both phases are shown in Fig. 1. For c-diamond, the room-temperature hole mobility () and electron mobility () are predicted to be 2543 and 2068 cm2V-1s-1, respectively, in good agreement with experimental values. The high mobilities of both carrier types overcome the limitation of most conventional semiconductors that only one carrier type (usually electrons) exhibits high mobility. Regarding h-diamond, its carrier mobilities are found to be much higher. At 300 K, the hole mobilities are = 5631 and = 5552 cm2V-1s-1, and the electron mobilities are = 11462 and = 28464 cm2V-1s-1, both of which are even superior to the typical values of most known high-hole-mobility semiconductors (e.g., Ge, with = 1820 cm2V-1s-1 [14]) and high-electron-mobility semiconductors (e.g., GaAs, with = 8900 cm2V-1s-1 [10]). While the predicted hole mobility agrees well with the result in Ref. 26, our calculated electron mobility significantly exceeds the reported value. We performed calculations using two different ab initio methods for cross-verification, and the results are in reasonable agreement [see Sec. S2 of Supplemental Material (SM) [[SeeSupplementalMaterialat][fordetailsoftheoreticalformulations, computationalmethods, high-tempreaturestabilityofh-diamond, effectsofsplit-offholebandandintervalleyscattering, andadditionalfigures.]sm] for details], confirming the exceptional charge transport property of h-diamond. Since h-diamond has also been predicted to have a room-temperature thermal conductivity of nearly 2000 Wm-1K-1 [4, 25], it may outperform other novel semiconductor materials like BAs, which has drawn much attention in recent years due to its high ambipolar mobility (about 1600 cm2V-1s-1) [24, 19] and high thermal conductivity (about 1300 Wm-1K-1) [11, 12, 20]. Additionally, our calculation shows that h-diamond has a wide bandgap (4.55 eV according to our calculation), and the molecular dynamics simulation at 1000 K indicates its high-temperature stability (see Sec. S3 in SM [1]). Therefore, h-diamond is expected as a promising candidate for applications in high-power and high-frequency electronics.
@@
To pave the way for the high-temperature application of h-diamond, the temperature dependence of carrier mobility is examined. Carrier mobility typically follows a power-law relationship with temperature, i.e., , if the scattering mechanism remains unchanged. However, as shown in Fig. 1(a), the temperature dependence of carrier mobility of h-diamond cannot be well described using the power law with a single exponent within 150–700 K, suggesting a transition in the predominant scattering mechanism. Since h-diamond is a nonpolar semiconductor, deformation potential scattering is expected to play a crucial role, which can be induced by either acoustic or optical phonons. In order to elucidate the contributions of different phonon modes to carrier scattering, we define a function of phonon energy as
where is the carrier mobility based on the self-energy time approximation [SERTA, see Eq. (S5) in SM [1]], with the scattering rate as
where is the transition rate from the initial state to the final state by the phonon . is analogous to , with the scattering rate replaced by
@@
where is the Gaussian function , with as the finite broadening. Consequently, is the carrier mobility that excludes scattering induced by phonons with an energy of , and thus a larger indicates a greater contribution of the corresponding phonons to limiting the mobility. As shown in Fig. 2, carrier scattering is mainly induced by phonons with an energy within two different ranges. The lower range is below 30 meV, indicating acoustic deformation potential (ADP) scattering. The higher range is between 138 and 160 meV, corresponding to optical modes. At 300 K, ADP is the predominant scattering mechanism. At higher temperatures, the relative contribution of optical phonons increases significantly, which dominate carrier-phonon scattering above about 400 K for holes and 450 K for electrons (see also Fig. S4 in SM [1]).
Given the same elemental composition and similar bonding characteristic of both phases, the reason for the much higher carrier mobility in h-diamond is of great interest. Mobility is usually associated with the carrier effective mass. A smaller effective mass typically indicates higher band velocities and a lower density of states (DOS), which give rise to faster carrier transport and reduced scattering phase space, respectively, thus leading to higher mobility. For example, it has been previously reported that the high hole mobility of c-diamond is mainly attributed to the small effective mass [23]. For comparison, we calculate the “mobility effective masses” [16] to quantify the effect of band curvature on carrier mobility, as listed in Table 1. And the DOS (per valley) is also evaluated for electronic states near band edges, as shown in Fig. S5 in SM [1]. Both quantities of h-diamond are comparable to those of c-diamond, supporting the small effective mass as an important contributor to the high mobility of h-diamond. However, it cannot explain the fact that the mobility in h-diamond is much higher than that in c-diamond, suggesting that other factors should play a crucial role. For example, one may attribute it to the differences in interband and intervalley scattering. For holes, the crystal-field splitting leads to a considerable energy gap (about 0.5 eV) between the valence band maximum (VBM) and the split-off hole (sh) band in h-diamond, which is likely to reduce the number of interband scattering channels. Regarding electrons, the conduction band minimum (CBM) of c-diamond at has a larger multiplicity compared with that of h-diamond (CBM at ), which might cause enhanced intervalley scattering. Nevertheless, our calculations indicate that both effects have only minor contributions (see Sec. S4 in SM [1] for details), which will thus not be included in further discussion.
@@
For holes in c-diamond, room-temperature mobility is predominantly limited by transverse acoustic (TA) phonons (see Fig. S6 in SM [1]), which agrees with previous study [23], and longitudinal acoustic (LA) phonons also play an appreciable role. These observations can be understood according to the selection rules (the notation for irreducible representations follows that of Ref. 5, see also Fig. S7 in SM [1]).
@@
Therefore, the symmetry places little restriction on acoustic modes. The scattering processes mediated by LA phonons are allowed for all these directions, and the TA modes are only partially forbidden.
@@
For h-diamond, the VBM is composed of and orbitals and has the symmetry, and the selection rules for in-plane scattering are
@@
which originate from the symmetry of the Lonsdaleite structure. Since both the initial and final states are odd under the (0001) mirror symmetry, the phonons that are odd under it, such as the out-of-plane TA mode, are forbidden from inducing such transitions, thereby reducing the number of scattering channels. For comparison, although both h-diamond and wurtzite structures exhibit the AB stacking and have hexagonal symmetry, the compound nature of the latter breaks the (0001) mirror symmetry, and thus all phonons can induce such in-plane scattering in terms of symmetry. And the selection rule for scattering parallel to the -axis is
@@
indicating that both TA modes ( symmetry) are forbidden. In conjunction with Eq. (7), it is found that the selection rules considerably suppress scattering in h-diamond caused by TA phonons, which predominantly limit hole mobility in c-diamond around room temperature. This could be a possible reason for the higher hole mobility in h-diamond.
@@
Meanwhile, h-diamond has four atoms in the primitive cell, thereby introducing six additional phonon modes. To elucidate its effect on hole mobility, we first focus on the “high-energy” ( meV) modes. In c-diamond, the transverse optical (TO) modes are allowed by selection rules [Eqs. (Exceptionally High Carrier Mobility in Hexagonal Diamond)–(6)] to induce scattering, while scattering mediated by longitudinal optical (LO) phonons is forbidden with the final states at the heavy-hole (hh) band 111It has been reported in Ref. 23 that the highest valence band along the direction exhibits a light-hole-like characteristic. To avoid ambiguity, we refer to the three top valence bands in order of descending energy as the heavy-hole, light-hole, and split-off hole bands throughout this work. along and directions. This also explains the trend in Fig. S6 that TO phonons become dominant at elevated temperatures, whereas LO phonons have a limited effect even at 700 K. In h-diamond, there are six optical branches within this range. Nevertheless, for both and directions, only two TO modes ( and symmetry) and one LO mode ( and symmetry) are allowed by the selection rules [Eqs. (7) and (8)]. Hence, the optical phonons in the high-energy region lead to few additional scattering channels in h-diamond, except that the LO-induced scattering is less restricted by selection rules, which could explain the trend in Fig. 1 that the relative difference in hole mobility between the two phases becomes smaller at higher temperatures.
@@
The “low-energy” ( meV) phonons in h-diamond along high-symmetry paths correspond to two interlayer shear modes and one interlayer breathing mode. For in-plane scattering, the shear vibrations are forbidden due to their odd parity under reflection across the (0001) plane [Eq. (7)]. Along the -direction, the breathing mode ( symmetry) is also forbidden [Eq. (8)]. In contrast, the breathing mode (for ) and the shear modes (for ) are allowed to induce scattering, leading to non-zero matrix elements. These modes are characterized by shear or stretching vibrations of bonds parallel to the -axis. As shown in Fig. 3, the and characteristic results in an in-plane distribution pattern of the charge density at VBM. Therefore, the spatial separation between the scattering potential and the charge distribution leads to weak carrier-phonon coupling in this regime.
@@
On the other hand, the in-plane distribution of the charge density at VBM also gives rise to anisotropy of the hole effective mass, which is larger parallel to the -axis. This is consistent with the calculated hole mobility based on SERTA, which is 23% lower along the direction than that along the directions at 300 K. Nonetheless, the Boltzmann transport equation (BTE) mobility reveals a different trend, which is almost identical along three Cartesian directions at 300 K, with the -component 24% higher than that of SERTA; at lower temperatures, the -component is even higher than the -components. This can be attributed to the weak scattering mediated by phonons with parallel to the -axis. According to Eq. (8), both TA modes cannot induce scattering. The LA phonons, albeit allowed to mediate scattering, correspond to stretching vibrations of bonds parallel to the -axis, thus resulting in suppressed scattering, similar to the case of interlayer shear and breathing modes discussed in the previous paragraph. In contrast, for , scattering by the LA and in-plane TA modes is allowed, both of which incorporate vibrations of bonds parallel to the basal plane, thereby giving rise to considerable carrier-phonon coupling. Hence, the -component of the band velocity of holes may be subject to little change due to scattering, leading to enhanced mobility along this direction, also explaining the underestimation of SERTA, which assumes complete loss of electronic momentum after each scattering process. At higher temperatures, the high-energy optical modes are activated, which incorporate in-plane opposite vibrations, and thus phonons with can also cause significant carrier-phonon coupling. Consequently, the hole mobility along the direction becomes lower than the -components above 300 K. At 700 K, the -component of BTE mobility is only 6% higher than its SERTA counterpart.

### v2

@@ Abstract
Hexagonal diamond (h-diamond), or Lonsdaleite, has been reported to be a wide-bandgap semiconductor with high thermal conductivity and hardness. Our ab initio calculations reveal its exceptionally high carrier mobility at room temperature. Along and directions, the hole mobilities are 5631 and 5552 cm2V-1s-1, and the electron mobilities are 11462 and 28464 cm2V-1s-1, respectively. These values are significantly superior to the mobility of most known semiconductors including cubic diamond. The small effective masses in h-diamond, comparable to those in cubic diamond, cannot explain its substantially higher mobility. Instead, two crucial mechanisms are uncovered: selection rules that considerably suppress hole scattering induced by transverse acoustic phonons, and a spatial decoupling effect where electronic wavefunctions concentrated in lattice interstitials lead to minimal overlap with scattering potentials.
I Introduction
@@
Diamond, long prized as a gemstone, is also regarded as the “ultimate” wide-bandgap semiconductor material because of its excellent physical properties [36]. While most semiconductors may show good performance in certain properties, they often face inherent limitations in other critical characteristics. For example, GaAs has high electron mobility yet low hole mobility and thermal conductivity [1]; -Ga2O3 has an ultrawide bandgap and large breakdown field, whereas its carrier mobility and thermal conductivity are unsatisfactory [7]. In contrast, diamond offers a unique combination of a series of key properties, including its high carrier mobility, high thermal conductivity, ultrawide bandgap, and high breakdown field [36]. These intrinsic characteristics position it at the frontier of future electronics operating under the most extreme conditions of high power, frequency, and temperature [31].
@@
In addition to the well-known cubic structure, diamond can also crystallise in the hexagonal phase. Hexagonal diamond (h-diamond), also referred to as Lonsdaleite [5], was both discovered in meteorites [13] and experimentally synthesised [2] in the 1960s. In contrast to cubic diamond (c-diamond) with a stacking sequence of ABCABC… along the directions, it exhibits an ABAB… stacking. Although far less explored than the well-known cubic phase, existing studies have shown that the properties of h-diamond are comparable to, or even superior to, those of c-diamond. For example, Pan et al. [23] revealed a 58% higher indentation strength of h-diamond than that of its cubic counterpart. Regarding lattice thermal conductivity, studies on h-diamond report scattered values [40, 30, 3], which are approximately 20–30% lower than that of c-diamond, yet still considerably higher than those of most known semiconductors.
@@
In this work, motivated by the recent success in the experimental synthesis of its highly-ordered bulk samples [37], we thoroughly investigate the carrier mobility of h-diamond using fully ab initio calculations, revealing that its mobility significantly exceeds that of most known semiconductor materials. Based on group theory and the spatial distributions of charge density and scattering potential, we further uncover two crucial mechanisms as the physical origins of its high mobility, thereby guiding the discovery and design of novel high-mobility materials.
II Results
II.1 Phonon-limited carrier mobility
@@
The calculated carrier mobilities of both phases are shown in Fig. 1. For c-diamond, the room-temperature hole mobility () and electron mobility () are predicted to be 2543 and 2068 cm2V-1s-1, respectively, in good agreement with experimental values [6]. Regarding h-diamond, its carrier mobilities are found to be much higher. At 300 K, the hole mobilities are = 5631 and = 5552 cm2V-1s-1. The electron mobilities are even higher, with = 11462 and = 28464 cm2V-1s-1. These values are superior to the typical mobility of most known high-mobility semiconductors, for example, Ge ( = 1820 cm2V-1s-1 [22]) for holes and GaAs ( = 8900 cm2V-1s-1 [14]) for electrons, also overcoming the limitation of most conventional semiconductors where only one carrier type exhibits high mobility. While the predicted hole mobility agrees well with the result in Ref. 41, our calculated electron mobility significantly exceeds the reported value. We carried out cross-verification using another ab initio method, and the results are in reasonable agreement with the data presented above [see Table S1 of the Supplementary Information (SI)], confirming the exceptional charge transport properties of h-diamond. Since h-diamond has also been predicted to have a room-temperature thermal conductivity of nearly 2000 Wm-1K-1 [3, 40], it may outperform other novel semiconductors such as BAs, which has attracted considerable attention in recent years due to its high ambipolar mobility (about 1600 cm2V-1s-1) [39, 32] and high thermal conductivity (about 1300 Wm-1K-1) [15, 18, 34]. Additionally, our calculation yields a wide bandgap of 4.55 eV for h-diamond, and a molecular dynamics (MD) simulation at 1000 K indicates its thermal stability (see Fig. S1 of the SI). These combined properties suggest h-diamond as a promising candidate for applications in high-power and high-frequency electronics.
@@
To pave the way for the high-temperature application of h-diamond, we examine the temperature dependence of its carrier mobility. Carrier mobility typically follows a power-law relationship with temperature, i.e., , if the scattering mechanism remains unchanged. However, as shown in Fig. 1(a), the temperature dependence of carrier mobility in h-diamond cannot be well described by a single power-law exponent over the range of 150–700 K, suggesting a transition in the dominant scattering mechanism, which will be analysed in the subsequent section.
II.2 Carrier scattering mechanisms
@@
Carrier mobility is governed by scattering mechanisms, an in-depth analysis of which can provide a better understanding of carrier transport behaviour at the microscopic level. In terms of electron-phonon (e-ph) interaction, since h-diamond is a non-polar semiconductor, deformation potential scattering is expected to play a central role, which can be induced by either acoustic or optical phonons. In order to elucidate the contributions of different phonon modes to carrier scattering, we define as a function of phonon energy [Eq. (S6) in SI]. A larger value of indicates a greater reduction in carrier mobility due to phonons with an energy around . As shown in Fig. 2, carrier scattering is primarily induced by phonons in two energy ranges. The lower range is below 30 meV, attributed to acoustic deformation potential (ADP) scattering. The higher range is between 138 and 160 meV, corresponding to optical modes. At lower temperatures around 300 K, ADP is the predominant scattering mechanism. Nevertheless, as shown in Figs. S2 and S3 of the SI, the relative contribution of the optical deformation potential (ODP) scattering increases significantly at higher temperatures, dominating carrier-phonon scattering above about 400 K for holes and 450 K for electrons.
Both scattering mechanisms considerably depend on the carrier effective mass. A smaller effective mass typically indicates higher band velocities and a lower density of states (DOS), which give rise to faster carrier transport and reduced scattering phase space, respectively, thereby increasing mobility. For example, the high hole mobility in c-diamond is attributed primarily to its small effective mass [38]. For comparison, we calculate the “mobility effective masses” [27] to quantify the effect of band curvature on carrier mobility, as listed in Table 1. We also evaluate the DOS (per valley) near band edges, as shown in Fig. S4 of the SI. Both quantities of h-diamond are comparable to those of c-diamond, indicating that the small effective mass contributes considerably to the high mobility in h-diamond. However, it cannot further explain the fact that the mobility in h-diamond is significantly higher.
@@
According to the relationship between the initial and final electronic states, carrier scattering can be divided into intraband and interband scattering, or intravalley and intervalley scattering. While intraband and intravalley scattering can always occur, the contributions of interband and intervalley scattering are primarily determined by the band structure. As shown in Fig. 3, the heavy-hole (hh) and light-hole (lh) bands (referred to as the highest and second highest valence bands, respectively) are close in energy, leading to interband scattering of holes. Regarding the conduction bands, the conduction band minimum (CBM) is away from the Brillouin zone centre, resulting in intervalley scattering between equivalent valleys.
For holes, the crystal-field splitting in h-diamond opens an appreciable energy gap between the valence band maximum (VBM) and the split-off hole (sh) band, whereas in c-diamond, this gap is only 16 meV at the point. It is thus natural to consider whether the lower hole mobility in c-diamond arises from the additional interband scattering channels that involve the sh band. However, the DOS near the VBM is actually comparable in both phases, as shown in Fig. S4 of the SI. To provide a quantitative analysis, we computed the room-temperature hole mobility with the sh band artificially excluded from the calculation. The result shows a slight increase in hole mobility of approximately 2%.
Regarding electrons, the CBM of c-diamond lies at the point and has a larger multiplicity than the CBM of h-diamond (located at ). To compare the relative contributions of intervalley scattering in both materials, we calculated the room-temperature electron mobility that only involves the scattering induced by phonons with wave vector lengths below a specific cutoff, i.e., . Since intravalley scattering is mediated by phonons with small , and the CBM valleys are well separated from their equivalent counterparts in both phases, the mobility- relationship directly decomposes the contributions of intervalley and intravalley scattering. As shown in Fig. S5 of the SI, intravalley scattering significantly dominates in h-diamond. In c-diamond, despite the multiple intervalley scattering paths (including -processes and -processes), the -processes play a negligible role, and the overall contribution of intervalley scattering is comparable to that in h-diamond, which is approximately 10%.
Therefore, the mobility difference between the two phases cannot be explained by either the interband scattering that involves the sh band or the intervalley scattering of electrons, suggesting that other factors should play a crucial role. According to the results above, we will henceforth focus on intravalley scattering and restrict our analysis to the hh and lh bands.
II.3 Origin of the higher hole mobility in h-diamond
@@
For holes in c-diamond, the room-temperature mobility is predominantly limited by transverse acoustic (TA) phonons (see Fig. S6 of the SI), which agrees with the previous study [38], and longitudinal acoustic (LA) phonons also play an appreciable role. These observations can be understood according to the selection rules (the notation for irreducible representations follows that of Ref. 4; see also Fig. S7 of the SI).
@@
Therefore, the symmetry imposes little restriction on acoustic modes. The scattering processes mediated by LA phonons are allowed for all these directions, and those induced by TA modes are only partially forbidden.
@@
For h-diamond, the VBM is composed of and orbitals and has symmetry. The selection rules for in-plane scattering are
@@
which originate from the symmetry of the Lonsdaleite structure. Since both the initial and final states are odd under (0001) mirror symmetry, phonons that are odd under it, such as the out-of-plane TA mode, are forbidden from inducing such transitions, thereby reducing the number of available scattering channels. By contrast, although both h-diamond and wurtzite structures exhibit the AB stacking and share hexagonal symmetry, the compound nature of the latter breaks the (0001) mirror symmetry. Consequently, all phonon modes are allowed by symmetry to induce in-plane scattering in wurtzite structures. The selection rule for scattering with is
@@
indicating that both TA modes ( symmetry) are forbidden. In conjunction with Eq. (4), the selection rules considerably suppress TA-scattering in h-diamond, which predominantly limits hole mobility in c-diamond around room temperature. This suppression significantly contributes to the higher hole mobility in h-diamond.
@@
Meanwhile, h-diamond contains four atoms in its primitive cell, giving rise to six additional phonon modes compared with c-diamond. To elucidate the effect of these extra modes on hole mobility, we first focus on the “high-energy” optical phonons with meV. In c-diamond, transverse optical (TO) modes are allowed by selection rules [Eqs. (II.3)–(II.3)] to induce hole scattering, while scattering mediated by longitudinal optical (LO) phonons is forbidden for final states at the hh band along and directions. This also explains the trend observed in Fig. S6 of the SI that TO phonons become dominant at elevated temperatures, whereas LO phonons have a limited effect even at 700 K. In h-diamond, there are six optical branches within this range. Nevertheless, for both and directions, only two TO modes ( and symmetry) and one LO mode ( and symmetry) are allowed by selection rules [Eqs. (4) and (5)] to mediate hole scattering. Hence, the high-energy optical phonons introduce few additional scattering channels in h-diamond. However, LO-induced scattering is less restricted by selection rules compared with c-diamond, which could account for the trend in Fig. 1 that the relative difference in hole mobility between the two phases becomes smaller at higher temperatures.
@@
The “low-energy” ( meV) optical phonons in h-diamond along high-symmetry paths consist of two interlayer shear modes and one interlayer breathing mode. For in-plane scattering, the shear modes are forbidden due to their odd parity under reflection across the (0001) plane. For hole scattering along the -axis, the breathing mode ( symmetry) is also forbidden [Eq. (5)]. In contrast, the breathing mode with and the shear modes with are allowed to mediate scattering, leading to non-zero matrix elements. These modes are characterised by shear or stretching vibrations of bonds parallel to the -axis. As shown in Fig. 4, the VBM exhibits an in-plane charge density distribution due to its - character. Therefore, the spatial separation between the in-plane charge distribution and the scattering potentials leads to weak carrier-phonon coupling in this regime.
@@
On the other hand, the in-plane charge density distribution at the VBM also leads to anisotropy in the hole effective mass, which is larger parallel to the -axis (Table 1). This is consistent with the hole mobility based on the self-energy relaxation time approximation (SERTA), which is 23% lower along than in the directions at 300 K. Nonetheless, the Boltzmann transport equation (BTE) reveals a different trend: at 300 K, the mobility is almost identical along three Cartesian directions; at lower temperatures, the -component is even higher than the -components. This enhancement of hole mobility along can be attributed to weak scattering by phonons with . According to Eq. (5), both TA modes are forbidden from inducing out-of-plane scattering. Although LA phonons are allowed to mediate scattering, they correspond to displacements of basal planes along the -axis, resulting in suppressed interactions with the in-plane charge distribution, similar to the case of interlayer shear and breathing modes discussed in the previous paragraph. In contrast, for , scattering by the LA and in-plane TA modes is allowed, both of which involve intralayer vibrations, leading to considerable carrier-phonon coupling. Hence, the -component of hole band velocity is subject to little change due to scattering, enhancing mobility along this direction. This also explains the underestimation of SERTA, which assumes a complete loss of carrier momentum after each scattering event and is likely to overestimate anisotropic scattering. At higher temperatures, high-energy optical modes exhibiting in-plane opposite vibrations are activated. As a result, phonons with can now induce significant scattering. Consequently, the hole mobility along the direction becomes lower than the -components above 300 K. At 700 K, the -component of BTE mobility is only 6% higher than its SERTA counterpart.
II.4 Origin of the higher electron mobility in h-diamond